# Optimal Quantization on Spherical Surfaces: Continuous and Discrete Models – A Beginner-Friendly Expository Study

### v1

@@ Abstract.
This expository paper provides a unified and pedagogical introduction to optimal quantization for probability measures supported on spherical surfaces, emphasizing both continuous and discrete settings. We first present a detailed geometric and analytical foundation of quantization on the unit sphere, including definitions of great and small circles, spherical triangles, geodesic distance, Slerp interpolation, the Fréchet mean, spherical Voronoi regions, centroid conditions, and quantization dimensions. Building upon this framework, we develop explicit continuous and discrete quantization models on spherical curves—great circles, small circles, and great circular arcs—supported by rigorous derivations and pedagogical exposition. For uniform continuous distributions, we compute optimal sets of $n$-means and the associated quantization errors on these curves; for discrete distributions, we analyze antipodal, equatorial, tetrahedral, and finite uniform configurations, illustrating convergence to the continuous model. The central conclusion is that for a uniform probability distribution supported on a one-dimensional geodesic subset of total length $L$, the optimal $n$-means form a uniform partition and the quantization error satisfies $V_{n}=L^{2}/(12n^{2})$. The exposition emphasizes geometric intuition, detailed derivations, and clear step-by-step reasoning, making it accessible to beginning graduate students and researchers entering the study of quantization on manifolds.
@@ 1. Introduction and Geometric Preliminaries
The present article is intended as a beginner-friendly expository companion to the author’s forthcoming comprehensive monograph that will provide a deeper and systematic treatment of quantization on spherical surfaces [12].
This section provides the essential background for the study of quantization on spherical surfaces, summarizing the geometric and analytical components that will be used throughout the paper.
@@ 1.1. Great Circle, Small Circle, and Spherical Triangle
1.2. Equator and the Prime Meridean, Latitude and Longitude.
@@ 1.4. Geodesic Distance via Geographical Coordinates
$\mathbf{x}_{1}\cdot\mathbf{x}_{2}$ $=(\cos\phi_{1}\cos\theta_{1})(\cos\phi_{2}\cos\theta_{2})+(\cos\phi_{1}\sin\theta_{1})(\cos\phi_{2}\sin\theta_{2})+(\sin\phi_{1})(\sin\phi_{2})$
$=\cos\phi_{1}\cos\phi_{2}\bigl(\cos\theta_{1}\cos\theta_{2}+\sin\theta_{1}\sin\theta_{2}\bigr)+\sin\phi_{1}\sin\phi_{2}$
$=\cos\phi_{1}\cos\phi_{2}\cos(\theta_{1}-\theta_{2})+\sin\phi_{1}\sin\phi_{2}.$
@@
$\Theta=\cos^{-1}\Big(\sin\phi_{1}\sin\phi_{2}+\cos\phi_{1}\cos\phi_{2}\cos(\theta_{1}-\theta_{2})\Big).$
@@
which is known as the geodesic distance between $P_{1}$ and $P_{2}$ via geographical coordinates.
@@ 1.6. The Two-Argument Arctangent Function $\operatorname{atan2}(y,x)$
The ordinary arctangent $\arctan(y/x)$ returns an angle only in the range $(-\tfrac{\pi}{2},\tfrac{\pi}{2})$, and therefore cannot distinguish between points lying in different quadrants of the Cartesian plane. To obtain the correct signed angle for any point $(x,y)\in\mathbb{R}^{2}\setminus\{(0,0)\}$, the two-argument arctangent function
$\theta=\text{atan2}(y,x)$
is used. Let $r=\sqrt{x^{2}+y^{2}}$. Then $\theta=\operatorname{atan2}(y,x)$ is defined as the unique angle in $(-\pi,\pi]$ satisfying
$\cos\theta=\frac{x}{r},\qquad\sin\theta=\frac{y}{r}.$
In other words, the function $\operatorname{atan2}(y,x)$ checks the signs of $x$ and $y$ to assign $\theta$ to the correct quadrant:
$\operatorname{atan2}(y,x)=\begin{cases}\arctan\!\left(\dfrac{y}{x}\right),&x>0,\\[10.00002pt] \arctan\!\left(\dfrac{y}{x}\right)+\pi,&y\geq 0,\;x<0,\\[10.00002pt] \arctan\!\left(\dfrac{y}{x}\right)-\pi,&y<0,\;x<0,\\[10.00002pt] +\tfrac{\pi}{2},&y>0,\;x=0,\\[3.99994pt] -\tfrac{\pi}{2},&y<0,\;x=0.\end{cases}$
@@ 1.7. Spherical Linear Interpolation (Slerp)
$a^{2}+b^{2}+2ab\cos s=\frac{\sin^{2}s}{\sin^{2}s}=1\text{ yielding }\|\gamma_{AB}(\tau)\|^{2},=\rho^{2},\text{ i.e., }\text{ i.e., }\|\gamma_{AB}(\tau)\|=\rho.$
@@
$e_{2}=\frac{\hat{u}_{B}-(\hat{u}_{A}\cdot\hat{u}_{B})\hat{u}_{A}}{\|\hat{u}_{B}-(\hat{u}_{A}\cdot\hat{u}_{B})\hat{u}_{A}\|}=\frac{\hat{u}_{B}-\cos s\,\hat{u}_{A}}{\sin\sigma}\text{ implying }\hat{u}_{B}=\cos s\,e_{1}+\sin s\,e_{2}.$
@@ 1.9. Quantization Error on the Sphere
Let $P$ be a Borel probability measure on a sphere $\mathbb{S}_{\rho}^{2}$ of radius $\rho$ equipped with the geodesic metric $d_{G}$. Let $\alpha=\{a_{1},a_{2},\dots,a_{n}\}\subset\mathbb{S}_{\rho}^{2}$ be a finite set of points. Such a set $\alpha$ is also refereed to as codebook and the elements as codepoints. Let $r\in\mathbb{R}$ with $r>0$. Then, the distortion error for $P$, of order $r>0$ associated with $\alpha$, denoted by $V_{r}(P;\alpha)$, is defined by
@@
$V_{n,r}(P)$ is called the $n$th quantization error of order $r$ for the probability measure $P$. When $r=2$, the problem corresponds to minimizing the mean squared geodesic distance between a random vector $X$ with distribution $P$ and its nearest codepoint. Then, $V_{n,r}(P)$ is then called the $n$th quantization error for $P$ with respect to the squared geodesic distance.
@@ 1.11. Optimal Quantizers on the Sphere
1.12. Centroid (Normalized Conditional Expectation) Condition (General Radius $\rho>0$)
@@
$\mathbb{S}^{2}_{\rho}=\{x\in\mathbb{R}^{3}:\|x\|=\rho\},$
and let $\alpha^{*}=\{a_{1}^{*},a_{2}^{*},\ldots,a_{n}^{*}\}\subset\mathbb{S}^{2}_{\rho}$ be an optimal set of $n$-means. Each representative point $a_{i}^{*}$ minimizes the local distortion functional
$\int_{R(a_{i}^{*}|\alpha^{*})}\|x-a_{i}\|^{2}\,dP(x)$
subject to the spherical constraint $\|a_{i}\|=\rho$.
Define the Euclidean conditional expectation (the unnormalized centroid) of the Voronoi region $R(a_{i}^{*}|\alpha^{*})$ by
$m_{i}:=\frac{\displaystyle\int_{R(a_{i}^{*}|\alpha^{*})}x\,dP(x)}{\displaystyle P(R(a_{i}^{*}|\alpha^{*}))}=\mathbb{E}[X\mid X\in R(a_{i}^{*}|\alpha^{*})].$
Since $\|m_{i}\|<\rho$ in general, the point $m_{i}$ lies inside the ball of radius $\rho$. To project it back onto the spherical surface $\mathbb{S}^{2}_{\rho}$, we normalize by $\rho$:
$a_{i}^{*}=\rho\,\frac{m_{i}}{\|m_{i}\|}=\rho\,\frac{\displaystyle\int_{R(a_{i}^{*}|\alpha^{*})}x\,dP(x)}{\displaystyle\Big\|\int_{R(a_{i}^{*}|\alpha^{*})}x\,dP(x)\Big\|}.$ (5)
Remark 1.12.1. The vector
$\int_{R(a_{i}^{*}|\alpha^{*})}x\,dP(x)$
represents the average Euclidean direction of points within the Voronoi region $R(a_{i}^{*}|\alpha^{*})$. Multiplying by $\rho/\|m_{i}\|$ rescales this average to lie exactly on the sphere of radius $\rho$, ensuring $a_{i}^{*}\in\mathbb{S}^{2}_{\rho}$. Geometrically, $a_{i}^{*}$ coincides with the intrinsic (Fréchet) mean of the conditional distribution $P(\,\cdot\,|\,R(a_{i}^{*}|\alpha^{*}))$ with respect to the geodesic distance $d_{G}(x,y)=\rho\,\arccos(\langle x,y\rangle/\rho^{2})$. Consequently, the collection
$\alpha^{*}=\{a_{i}^{*}:1\leq i\leq n\}$
forms a spherical optimal set of $n$-means for the probability distribution $P$ on $\mathbb{S}^{2}_{\rho}$.
@@ 1.13. Applications
  • Directional statistics and meteorology (e.g., wind directions, orientations).
@@ 2.1. Coordinates and metric
where $\phi\in[0,\pi]$ is the latitude and $\theta\in[0,2\pi)$ is the longitude of the point.
@@
When $x,y$ lie on the same great circle, $d_{G}(x,y)$ is simply their minimal angular separation.

### v2

# To appear, Mathematics

@@ Abstract.
This expository paper provides a unified and pedagogical introduction to optimal quantization for probability measures supported on spherical curves and discrete subsets of the sphere, emphasizing both continuous and discrete settings. We first present a detailed geometric and analytical foundation for intrinsic quantization on the unit sphere, including definitions of great and small circles, spherical triangles, geodesic distance, Slerp interpolation, the Fréchet mean, spherical Voronoi regions, centroid conditions, and quantization dimensions.
Building upon this framework, we develop explicit continuous and discrete quantization models on spherical curves, namely great circles, small circles, and great circular arcs— supported by rigorous derivations and pedagogical exposition. For uniform continuous distributions, we compute optimal sets of $n$-means and the associated quantization errors on these curves; for discrete distributions, we analyze antipodal, equatorial, tetrahedral, and finite uniform configurations, illustrating convergence to the continuous model.
The central conclusion is that for a uniform probability distribution supported on a one-dimensional geodesic subset of total length $L$, the optimal $n$-means form a uniform partition and the quantization error satisfies $V_{n}=L^{2}/(12n^{2})$. The exposition emphasizes geometric intuition, detailed derivations, and clear step-by-step reasoning, making it accessible to beginning graduate students and researchers entering the study of quantization on manifolds. This article is intended as an expository and tutorial contribution, with the main emphasis on geometric reformulation and pedagogical clarity of intrinsic quantization on spherical curves, rather than on the development of new asymptotic quantization theory.
@@ 1. Introduction and Geometric Preliminaries
While the general problem of optimal quantization on spherical surfaces is inherently high-dimensional and analytically challenging, the focus of this paper is deliberately more specific and pedagogical. We concentrate on intrinsic quantization problems for probability measures supported on one-dimensional spherical curves, namely great circles, small circles, and geodesic arcs. In these settings, the sphere induces a one-dimensional Riemannian structure via arc-length, allowing the intrinsic quantization problem to reduce to a transparent and fully explicit one-dimensional model.
The central theme of the paper is that, once the correct intrinsic metric is adopted, quantization on such spherical curves reduces to classical one-dimensional quantization with respect to arc-length. This reduction explains why the optimal Voronoi cells are contiguous intervals and why, under uniform density, the optimal codepoints are the midpoints of these intervals.
@@
Before entering the detailed geometric and analytic developments, we summarize the guiding principle underlying the main results of this paper. While optimal quantization on spherical surfaces is, in general, a high-dimensional and analytically challenging problem, the focus of this paper is deliberately more specific. We concentrate on intrinsic quantization problems for probability measures supported on one-dimensional spherical curves, namely great circles, small circles, and geodesic arcs. Although the ambient space is the two-dimensional sphere, once the support of the probability measure is constrained to such a spherical curve, the intrinsic quantization problem effectively reduces to a one-dimensional Riemannian model with arc-length parameterization. This reduction principle allows for a transparent and explicit analysis.
Nature and contribution of the paper. The results presented here recover, in the intrinsic spherical setting, the classical one-dimensional uniform quantization formula
$V_{n}=\frac{L^{2}}{12n^{2}},$
where $L$ denotes the intrinsic length of the support curve. Accordingly, the primary contribution of this paper is not the discovery of new quantization asymptotics, but rather a geometric re-framing of known one-dimensional results within the intrinsic geometry of the sphere. The paper is intended as a tutorial and pedagogical note, highlighting how intrinsic metrics, geodesic Voronoi structures, and arc-length parameterization lead naturally to transparent and explicit solutions on spherical curves.
The reduction principle described above is summarized in the following conceptual theorem.
Theorem 1.1 (Conceptual reduction to the one-dimensional uniform case).
Let $C\subset\mathbb{S}^{2}_{\rho}$ be a smooth closed geodesic curve of total intrinsic length $L$ (e.g. a great circle or a small circle), and let $P$ be the uniform probability measure on $C$ with respect to arc-length. Then the intrinsic quantization problem on $(C,d_{G})$ with squared geodesic distortion is equivalent to the classical one-dimensional uniform quantization problem on a circle of length $L$.
In particular, for each $n\geq 1$, the optimal set of $n$-means consists of $n$ equally spaced points along $C$, each Voronoi region has length $L/n$, each codepoint is the geodesic midpoint of its cell, and the quantization error satisfies
$V_{n}=\frac{L^{2}}{12n^{2}}.$
The present article is intended as a beginner-friendly expository companion to the author’s forthcoming comprehensive monograph, which will provide a deeper and more systematic treatment of quantization on spherical surfaces [12]. Accordingly, we emphasize geometric intuition, explicit derivations, and pedagogical clarity throughout. This section provides the essential background for the study of quantization on spherical surfaces, summarizing the geometric and analytical components that will be used throughout the paper.
@@ 1.2. Great Circle, Small Circle, and Spherical Triangle
1.3. Equator and the Prime Meridian, Latitude and Longitude.
@@ 1.5. Geodesic Distance via Geographical Coordinates
$\mathbf{x}_{1}\cdot\mathbf{x}_{2}$ $=\rho^{2}\left[(\cos\phi_{1}\cos\theta_{1})(\cos\phi_{2}\cos\theta_{2})+(\cos\phi_{1}\sin\theta_{1})(\cos\phi_{2}\sin\theta_{2})+(\sin\phi_{1})(\sin\phi_{2})\right]$
$=\rho^{2}\left[\cos\phi_{1}\cos\phi_{2}\bigl(\cos\theta_{1}\cos\theta_{2}+\sin\theta_{1}\sin\theta_{2}\bigr)+\sin\phi_{1}\sin\phi_{2}\right]$
$=\rho^{2}\left[\cos\phi_{1}\cos\phi_{2}\cos(\theta_{1}-\theta_{2})+\sin\phi_{1}\sin\phi_{2}\right].$
@@
$\Theta=\cos^{-1}\Big(\frac{x_{1}\cdot x_{2}}{\rho^{2}}\Big)=\cos^{-1}\Big(\sin\phi_{1}\sin\phi_{2}+\cos\phi_{1}\cos\phi_{2}\cos(\theta_{1}-\theta_{2})\Big).$
@@
which is known as the geodesic distance between $P_{1}$ and $P_{2}$ via geographical coordinates on a sphere of radius $\rho$.
@@ 1.7. The two-argument arctangent function $\operatorname{atan2}(y,x)$
To represent angles in the full range $(-\pi,\pi]$ without ambiguity, we use the two-argument arctangent function $\theta=\operatorname{atan2}(y,x)$. For any $(x,y)\neq(0,0)$, this function returns the unique angle $\theta\in(-\pi,\pi]$ such that
$\cos\theta=\frac{x}{\sqrt{x^{2}+y^{2}}},\qquad\sin\theta=\frac{y}{\sqrt{x^{2}+y^{2}}}.$
Unlike the single-variable function $\arctan(y/x)$, the function $\operatorname{atan2}(y,x)$ automatically accounts for the correct quadrant of the point $(x,y)$ and is standard in geometry, navigation, and numerical implementations.
In this paper, $\operatorname{atan2}$ is used only to define longitude coordinates consistently and plays no further conceptual role in the quantization analysis.
@@ 1.8. Spherical Linear Interpolation (Slerp)
$a^{2}+b^{2}+2ab\cos s=\frac{\sin^{2}s}{\sin^{2}s}=1\text{ yielding }\|\gamma_{AB}(\tau)\|^{2}=\rho^{2},\text{ i.e., }\|\gamma_{AB}(\tau)\|=\rho.$
@@
$e_{2}=\frac{\hat{u}_{B}-(\hat{u}_{A}\cdot\hat{u}_{B})\hat{u}_{A}}{\|\hat{u}_{B}-(\hat{u}_{A}\cdot\hat{u}_{B})\hat{u}_{A}\|}=\frac{\hat{u}_{B}-\cos s\,\hat{u}_{A}}{\sin s}\text{ implying }\hat{u}_{B}=\cos s\,e_{1}+\sin s\,e_{2}.$
@@ 1.10. Quantization Error on the Sphere
Let $P$ be a Borel probability measure on a sphere $\mathbb{S}_{\rho}^{2}$ of radius $\rho$ equipped with the geodesic metric $d_{G}$. Let $\alpha=\{a_{1},a_{2},\dots,a_{n}\}\subset\mathbb{S}_{\rho}^{2}$ be a finite set of points. Such a set $\alpha$ is also referred to as codebook and the elements as codepoints. Let $r\in\mathbb{R}$ with $r>0$. Then, the distortion error for $P$, of order $r>0$ associated with $\alpha$, denoted by $V_{r}(P;\alpha)$, is defined by
@@
$V_{n,r}(P)$ is called the $n$th quantization error of order $r$ for the probability measure $P$. When $r=2$, the problem corresponds to minimizing the mean squared geodesic distance between a random vector $X$ with distribution $P$ and its nearest codepoint. Then, $V_{n,r}(P)$ is called the $n$th quantization error for $P$ with respect to the squared geodesic distance.
@@ 1.12. Optimal Quantizers on the Sphere
1.13. Extrinsic centroid condition and its relation to intrinsic (Fréchet) means
@@
$\mathbb{S}^{2}_{\rho}:=\{x\in\mathbb{R}^{3}:\|x\|=\rho\},$
equipped with the geodesic distance $d_{G}$. Let
$\alpha^{*}=\{a_{1}^{*},a_{2}^{*},\dots,a_{n}^{*}\}\subset\mathbb{S}^{2}_{\rho}$
be an optimal set of $n$-means for $P$ (with respect to squared geodesic distortion), and let $R(a_{i}^{*}\mid\alpha^{*})$ denote the associated spherical Voronoi regions.
Extrinsic centroid condition. In many applied and algorithmic treatments of quantization on spheres (notably spherical $k$-means), one considers an extrinsic formulation obtained by viewing $\mathbb{S}^{2}_{\rho}$ as embedded in $\mathbb{R}^{3}$ and minimizing Euclidean squared distance subject to the spherical constraint. More precisely, for each Voronoi region $R(a_{i}^{*}\mid\alpha^{*})$ one minimizes
$\int_{R(a_{i}^{*}\mid\alpha^{*})}\|x-a\|^{2}\,dP(x)\quad\text{subject to }\|a\|=\rho.$
Define the (unnormalized) Euclidean conditional expectation over the Voronoi region by
$m_{i}:=\int_{R(a_{i}^{*}\mid\alpha^{*})}x\,dP(x).$
Since typically $\|m_{i}\|<\rho$, the constrained minimizer is obtained by radial projection of $m_{i}$ onto $\mathbb{S}^{2}_{\rho}$, yielding the extrinsic centroid
$a_{i}^{*}\;=\;\rho\,\frac{m_{i}}{\|m_{i}\|}\;=\;\rho\,\frac{\displaystyle\int_{R(a_{i}^{*}\mid\alpha^{*})}x\,dP(x)}{\displaystyle\left\|\int_{R(a_{i}^{*}\mid\alpha^{*})}x\,dP(x)\right\|}.$ (5)
We emphasize that (5) is an extrinsic (embedded/chordal) centroid rule: it arises from Euclidean squared loss in $\mathbb{R}^{3}$ together with the constraint $\|a\|=\rho$.
Relation to intrinsic (Fréchet) means. The intrinsic (Fréchet) mean on $\mathbb{S}^{2}_{\rho}$ is defined as a minimizer of the geodesic Fréchet functional
$q\;\mapsto\;\int_{\mathbb{S}^{2}_{\rho}}d_{G}(x,q)^{2}\,dP(x).$
In general, the extrinsic centroid (5) does not coincide with the intrinsic Fréchet mean, since the intrinsic problem depends nonlinearly on the Riemannian geometry through the exponential and logarithm maps.
For a Voronoi region $R\subset\mathbb{S}^{2}_{\rho}$, the correct intrinsic first-order optimality condition for a minimizer $a\in\mathbb{S}^{2}_{\rho}$ of
$F_{R}(a):=\int_{R}d_{G}(x,a)^{2}\,dP(x)$
is the vanishing of the Riemannian gradient, which can be expressed (when $R$ lies in a domain where the logarithm map is well-defined) as
$\int_{R}\log_{a}(x)\,dP(x)=0,$ (6)
where $\log_{a}(\cdot)$ denotes the Riemannian logarithm map at $a$ on the sphere. In general, (6) is an implicit condition and does not yield a closed-form centroid formula.
Coincidence in the one-dimensional uniform models studied here. Although intrinsic and extrinsic centroids differ in general, they coincide in several important situations, including:
  • when the support of the conditional distribution $P(\cdot\mid R(a_{i}^{*}\mid\alpha^{*}))$ is contained in a sufficiently small geodesic ball (so the intrinsic Fréchet functional is strictly convex), and
  • when the conditional distribution is symmetric with respect to a geodesic midpoint.
In particular, for the one-dimensional uniform models treated in Sections 4–7, each Voronoi region is a geodesic interval (arc), and the intrinsic Fréchet mean reduces to the geodesic midpoint of that interval. Consequently, in these settings the representative points produced by the extrinsic centroid rule (5) agree with the intrinsic minimizers for squared geodesic distortion.
Remark. A fully intrinsic “centroid condition” for optimal quantizers on $\mathbb{S}^{2}_{\rho}$ can be developed by analyzing the Riemannian gradient of the Fréchet functional and using the logarithm map, leading to (6). Since the focus of this paper is on explicit and pedagogical models where intrinsic means are available in closed form (and coincide with midpoints on geodesic intervals), we restrict attention to the clarification above.
@@ 1.15. Applications
  • Directional statistics, and earth and planetary sciences (e.g., wind directions, orientations).
@@ 2. Geometry of the Unit Sphere
Coordinate conventions used throughout the paper. We work on the sphere $S^{2}_{\rho}=\{x\in\mathbb{R}^{3}:\|x\|=\rho\}$ using standard spherical coordinates $(\rho,\theta,\phi)$ defined by $x=\rho(\sin\phi\cos\theta,\;\sin\phi\sin\theta,\;\cos\phi),$ where $\rho>0,\qquad\theta\in[0,2\pi),\qquad\phi\in[0,\pi].$ Here $\phi$ denotes the colatitude (measured from the north pole), while $\theta$ denotes the longitude. For readers familiar with geographical coordinates, the latitude $\lambda\in[-\pi/2,\pi/2]$ is related to $\phi$ by $\lambda=\frac{\pi}{2}-\phi.$ Throughout the paper, we consistently use the colatitude $\phi$ rather than latitude $\lambda$, unless explicitly stated otherwise.
Normalization to the unit sphere. Throughout this section we work on the unit sphere $\mathbb{S}^{2}\subset\mathbb{R}^{3}$. This choice is made purely for notational convenience and involves no loss of generality. Indeed, for a sphere $\mathbb{S}^{2}_{\rho}$ of arbitrary radius $\rho>0$, the intrinsic (geodesic) distance and arc-length scale linearly with $\rho$, i.e., $d_{G,\rho}(x,y)=\rho\,d_{G,1}(x/\rho,y/\rho)$. Consequently, any geodesic curve of intrinsic length $L$ on $\mathbb{S}^{2}_{\rho}$ corresponds, after normalization, to a curve of length $L/\rho$ on the unit sphere. Since the quantization problems studied in this paper depend only on the intrinsic one-dimensional geometry through arc-length, all results obtained on $\mathbb{S}^{2}$ extend immediately to $\mathbb{S}^{2}_{\rho}$ by the simple rescaling $L\mapsto\rho L$. Working on the unit sphere therefore allows for a cleaner exposition while preserving full generality.
@@ 2.1. Coordinates and metric
where $\phi\in[0,\pi]$ is the colatitude and $\theta\in[0,2\pi)$ is the longitude of the point.
@@
When $x,y$ lie on the same great circle, $d_{G}(x,y)$ is simply their minimal angular separation. When no ambiguity arises, we write $d_{G}$ for the geodesic distance on the underlying manifold.